OpenStudy (anonymous):

(x, 2.5) and (6, 4) If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. a.x = 10.1 c. x = 12.4 b .x = 9.6 d.x = 3.6

OpenStudy (ipwnbunnies):

It's saying what would 'x' equal with the equation is true: x1*y1 = x2*y2

OpenStudy (anonymous):

explain

OpenStudy (ipwnbunnies):

\[2.5x = 6*4 \]

OpenStudy (ipwnbunnies):

You're given two sets of coordinates, (X1, Y1) and (X2, Y2). Let's go with the first point. X1*Y1 = k, where 'k' is some random number. (X2,Y2) must also be equal to k. Therefore, X1*Y1 = X2*Y2

OpenStudy (anonymous):

If (x1, y1) and (x2, y2) satisfy xy = k, then x1y1 = x2y2. Find x or y as indicated. (4.5, y) and (12, 6) a. y = -2 c. y = 16 b. y = 0 d. y = 12

OpenStudy (anonymous):

can you help me with this one

OpenStudy (ipwnbunnies):

Make an equation, and solve for the variable.

OpenStudy (ipwnbunnies):

.-. It's the same type of problem. Read it. The two sets of coordinates still have to satisfy X1*Y1 = X2*Y2

OpenStudy (anonymous):

i still dont understand

OpenStudy (ipwnbunnies):

It's the same thing I explained in the first question.

OpenStudy (ipwnbunnies):

(X1, Y1) = (4.5, Y) (X2, Y2) = (12, 6) X1*Y1 = X2*Y2 Can you make an equation now, by substituting. :3

OpenStudy (anonymous):

The wattage rating of an appliance is given as W, in watts, and varies jointly as the resistance, R, in ohms, and as the square of the current, I, in amperes. If the wattage is 36 watts when the resistance is 100 ohms and the current is 0.6 amperes, find the wattage when the resistance is 300 ohms and the current is 0.5 amperes.